If all real values of x obtained from the equation
4^x – (a – 3)2^x +a – 4 = 0 are non – positive then a lies in 
a)(4,5]
b)(0,4)
c)(4,infinity)
d)none of these

Grade:

1 Answers

Arun
5 years ago
let 2^x = t
 
then
 
t^2 – (a – 3)t + a – 4 = 0
 
t^2  + t – (a – 4)t + a – 4 = 0
 
t (t + 1) – (a – 4) (t + 1) = 0
 
t = – 1
 
and 
t = a – 4
 
2^x = a – 4
 
0
 
4